OpenStudy (jazzyfa30):

Solve the equation by factoring: x2 - 81 = 0

hartnn (hartnn):

81 is the square of which number?

OpenStudy (jazzyfa30):

9

hartnn (hartnn):

right! so, we have \(x^2-9^2 =0\) now use \(a^2-b^2=(a+b)(a-b)\) to factor that!, what u get ?

hartnn (hartnn):

a=x, b=9

OpenStudy (jazzyfa30):

-9 or 9

OpenStudy (pradius):

(x+9)(x-9)

hartnn (hartnn):

correct! good :)

OpenStudy (jazzyfa30):

ok